OpenStudy (dmezzullo):

Imagine you are teaching a fellow student how to solve: 4 − (x + 2) = 6 − 5(x + 4) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (dmezzullo):

@sirm3d

OpenStudy (dmezzullo):

@mikala1

OpenStudy (dmezzullo):

@sirm3d

OpenStudy (anonymous):

here 4 - (x +2) = 6 - 5 (x+4) --multiply through to get rid of the parenthesis 4 -x -2 = 6 -5x - 20 -- combine like terms 2 - x = -14 - 5x -- get the x's on one side -x + 5x = -14 - 2 4x = -16 x = -4 check... 4 -x - 2 = 6 - 5x - 20 4 -(-4) -2 = 6 -5(-4) - 20 4 + 4 - 2 = 6 + 20 - 20 8 - 2 = 6 6 = 6 (correct)

OpenStudy (dmezzullo):

tyvm mik

OpenStudy (anonymous):

welcome

OpenStudy (dmezzullo):

medal 4 me?